OpenStudy (anonymous):

Solve the integral: sin^3(x) * cos^5(x) Which trigonometric rule(s) do you use and how?

OpenStudy (anonymous):

from looking at this i might actually lop off one of the sine's

OpenStudy (anonymous):

maybe write it like this: sin^2(x)cos^5(x)sin(x), then i am sure you can use the Pythagorean identity

OpenStudy (anonymous):

Substitute \(t = \sin^2x\) \[dx = \frac{dt}{2\sin x \cos x}\]

OpenStudy (anonymous):

\[\int \frac{t* \cancel{\cos x * \sin x } (\cos^2 x)^2}{2 \cancel{\sin x * \cos x}}dt\]

OpenStudy (anonymous):

\[\int \frac{t ( 1 - t)^2}{2}dt\]

OpenStudy (anonymous):

\[\frac{1}{2}* ( \int (t^3 + t - 2t^2)dt)\]

OpenStudy (amistre64):

simpler is always better :) unless you need the practice

OpenStudy (anonymous):

Indeed :)

OpenStudy (anonymous):

there are many ways to skin a cat

OpenStudy (anonymous):

;) Thanks